### v2

@@ Abstract
Averaging checkpoints along the training trajectory is a simple yet powerful approach to improve the generalization performance of Machine Learning models and reduce training time. Motivated by these potential gains, and in an effort to fairly and thoroughly benchmark this technique, we present an extensive evaluation of averaging techniques in modern Deep Learning, which we perform using AlgoPerf (Dahl et al., 2023), a large-scale benchmark for optimization algorithms. We investigate whether weight averaging can reduce training time, improve generalization, and replace learning rate decay, as suggested by recent literature. Our evaluation across seven architectures and datasets reveals that averaging significantly accelerates training and yields considerable efficiency gains across all considered workloads, at the price of a minimal implementation and memory cost, while mildly improving generalization. Finally, we explore the relationship between averaging and learning rate annealing and show that combining the two achieves optimal performance.
@@ 1 Introduction
Training Deep Learning models is both resource-expensive and time-consuming. A principled and simple approach to speed up training, dating back to Polyak (1990) and Ruppert (1988), involves averaging the model’s weights across training iterations. This can either be performed online during training or post hoc by averaging checkpoints, effectively creating an ensemble of models at minimal additional cost. Previous studies have shown that weight averaging (WA) techniques can improve generalization (Merity et al., 2017; Gupta et al., 2020; Kaddour, 2022; Melis, 2023), increase robustness (Morales-Brotons et al., 2024), smooth loss landscapes (Izmailov et al., 2019), and accelerate convergence (Athiwaratkun et al., 2018; Li et al., 2023; Sanyal et al., 2023). Recent studies have also explored the connection between learning rate decaying and weight averaging (Sandler et al., 2023; Hägele et al., 2024), and used the latter to develop schedule-free algorithms (Defazio et al., 2024b).
@@
In this work, we present the largest evaluation of averaging techniques in modern Deep Learning, which we perform using AlgoPerf (Dahl et al., 2023), a collection of large-scale workloads and architectures developed to provide a unified benchmark for optimization algorithms. Framing our analysis in this setting provides a carefully designed evaluation framework, enables comparisons against strong, heavily tuned baselines, and allows drawing broad and robust conclusions.
@@
    We show that averaging can significantly accelerate training dynamics across seven different models and architectures and estimate a 12% reduction in GPU-hours to train the entire AlgoPerf suite up to the validation target compared to our baseline. We find this effect to be consistent across hyperparameters and observe encouraging speed-ups even on a more sophisticated optimizer like Distributed Shampoo (Shi et al., 2023).
@@
    Finally, we demonstrate how averaging checkpoints can act as a proxy for a shorter learning rate decay, but that it cannot fully replace learning rate schedules, at least within the explored variants, addressing an important question about the role of these techniques (Hägele et al., 2024; Defazio et al., 2024b).
@@ 2 Related Work
The idea of averaging iterates along a stochastic optimization trajectory dates back to Polyak (1990) and Ruppert (1988), and is often referred to as Polyak–Ruppert averaging. It has been extensively studied in the stochastic approximation framework (Polyak & Juditsky, 1992; Bach & Moulines, 2013; Neu & Rosasco, 2018; Lakshminarayanan & Szepesvari, 2018), and is a common technique to derive convergence guarantees in both convex and nonconvex optimization (Garrigos & Gower, 2023).
@@
Weight averaging techniques have seen extensive use in the Deep Learning community, often without explicit recognition or emphasis, despite their effectiveness. Influential works such as Szegedy et al. (2016), Vaswani et al. (2017), and Merity et al. (2017) have demonstrated their ability to enhance model performance and mitigate overfitting.
@@
The work of Izmailov et al. (2019) sparked renewed interest in weight averaging by demonstrating how averaging points along the SGD trajectory leads to wider minima and improves generalization performance. Their approach, Stochastic Weight Averaging (SWA), has since been applied to semi-supervised learning (Tarvainen & Valpola, 2018; Athiwaratkun et al., 2018), low-precision training (Yang et al., 2019), domain generalization tasks (Cha et al., 2021; Arpit et al., 2022), and meta-optimization (Li et al., 2023).
@@
In the original formulation of SWA, a pretrained model is trained with a cyclical or constant learning rate, and multiple checkpoints are collected and later averaged. Kaddour (2022) proposed Latest Weight Averaging (LAWA), an online algorithm that averages the latest checkpoints in a rolling window, showing significant speed-ups on vision and language tasks. Further modifications of LAWA demonstrated notable boosts in pretraining modern decoder-only language models (Sanyal et al., 2023). A valuable alternative to moving window averaging techniques like LAWA is Exponential Moving Averaging (EMA) (Li et al., 2024; Morales-Brotons et al., 2024; Arpit et al., 2022). It retains similar advantages of rolling window averaging, and constitutes an indispensable technique for high-quality image synthesis models such as GANs and diffusion models (Yazıcı et al., 2019; Song et al., 2021; Karras et al., 2024).
@@
Classical stochastic smooth convex optimization rates showcase a tight link between WA and learning rate annealing, suggesting a practical interplay between these techniques (see e.g. Theorem 5.3. in Garrigos & Gower (2023)). Intuitively, averaging weights along the training trajectory reduces noise and might act as a proxy for learning rate decay. In fact, Sandler et al. (2023) proved the theoretical and empirical equivalence between WA and decaying learning rate for SGD. However, despite this appealing result, modern Deep Learning models are still predominantly trained with learning rate annealing (Hägele et al., 2024), even when maintaining an EMA of model weights (DeepSeekAI, 2025). A recent study by Defazio et al. (2024b) specifically investigates this connection, proposing an approach that fully replaces learning rate schedules with iterate averaging and demonstrating strong performance on the same benchmark used in this analysis. Whereas Defazio et al. (2024b) incorporates averaging directly into the optimization procedure, we explore a different flavor of averaging, where the averaged weights do not influence the updates—akin to Polyak averaging, SWA, LAWA, and EMA.
@@
Finally, a different but notable approach that leverages the benefits of averaging is model soups (Wortsman et al., 2022). In this case, multiple models are trained with different hyperparameter configurations and later aggregated, resulting in an ensemble with improved accuracy and robustness.
In this work, we demonstrate the benefits of weight averaging techniques on a challenging optimization benchmark (Dahl et al., 2023), hoping to encourage broader adoption of these methods in training large-scale Machine Learning models.
@@ 3 Experimental Setup
We explore two flavors of weight averaging discussed in Section 2: LAWA (Kaddour, 2022) and EMA. For both approaches, we update a buffer that stores previous information about the model history. In the first case, we update a circular queue of length $L$, which stores recent checkpoints, while in the latter case, we maintain an exponential moving average of the model parameters, with coefficient $\gamma$. We update the averaging buffer every $\nu$ iterations, and explore the effect of such hyperparameter together with the value of $L$ and $\gamma$. In most scenarios, we save checkpoints of the baseline algorithm and run averaging schemes offline, but when testing collecting consecutive checkpoints, we use online versions of LAWA and EMA to avoid excessive disk storage. We discuss implementation details of the algorithms, along with their practical implications and considerations, in Appendix E.
@@
We conduct our analysis on AlgoPerf (Dahl et al., 2023), a large suite of Deep Learning workloads, which provides a comprehensive benchmark for testing optimization algorithms. The benchmark is composed of eight workloads, each defined by a dataset, a model architecture, and a predefined target metric on a held-out set, designed to represent optimal performance on such a workload. We note that AlgoPerf has been developed following option (A) and that significant effort has been dedicated to deriving challenging target scores. Thus, it provides an ideal setting to evaluate the impact of weight averaging techniques applied to strong optimization algorithms and compare their performance against heavily tuned baselines. We also make use of the same datasets and architectures to score algorithms by means of their generalization capabilities (when following option (B)). We consider the following workloads from the AlgoPerf suite: (i) a DLRMsmall model on Criteo 1TB dataset for click-through rate prediction; (ii) U-Net on FastMRI for medical image reconstruction; (iii) ViT on ImageNet-1k for image classification; (iv) a GNN model on OGBG for graph-based prediction; (v) a Transformer-Big on WMT for machine translation; (vi) a Conformer for speech recognition; (vii) a DeepSpeech model on LibriSpeech for speech-to-text. We exclude ImageNet‐ResNet from this analysis because, as shown in Dahl et al. (2023), no optimization algorithm successfully solves this task alongside the other workloads, leaving no baseline on which to build. We refer to Appendix A for more details on the workloads.
@@
We build on top of existing optimization algorithms that perform well on AlgoPerf, enhancing them with LAWA or EMA. Unless otherwise specified, we use NadamW (Timothy, 2016; Loshchilov & Hutter, 2017a) as our baseline and name the resulting algorithms NadamW + LAWA and NadamW + EMA, respectively. Additionally, we investigate the potential benefits of combining a higher-order optimizer, such as Distributed Shampoo (Shi et al., 2023), with these averaging methods, exploring how this approach might further improve training efficiency. Given the high computational cost of the benchmark, we conduct most of the analysis using the best performing hyperparameters in Dahl et al. (2023) and Kasimbeg et al. (2025): this gives us strong baselines, avoiding the burden of expensive hyperparameter tuning. Additionally, we ablate on the role of the learning rate in each analysis and study in more detail the consequences of changing the learning rate schedule. Given this reference algorithm, we add LAWA or EMA on top of it, tuning only the hyperparameters of the averaging scheme and leaving the other baseline hyperparameters fixed, including learning rate schedule and weight decay. Unless otherwise specified, or when explicitly ablating on it, we train using a cosine learning rate schedule (Loshchilov & Hutter, 2017b).
@@
The impact of an averaging scheme largely depends on the horizon over which it is applied and on the frequency at which checkpoints are collected. A long horizon may prioritize outdated model checkpoints and generalize worse, whereas a short one may be suboptimal or result in a serious computational overhead. At the same time, updating the averaging buffers every step is only possible in the online version of LAWA and EMA, and collecting checkpoints too rarely might reduce the benefits of averaging. Previous works have collected checkpoints at each training step (Morales-Brotons et al., 2024; Hägele et al., 2024), or the end of each epoch (Kaddour, 2022), and investigated the optimal horizon and update frequency on a single task (Sanyal et al., 2023). We explore combinations and interactions of these variables across architectures and objectives, and compare LAWA with EMA.
@@ 4 Speeding up Training
We ask whether equipping a strong baseline with LAWA or EMA can reduce the number of steps required to reach a predefined validation target (option A in Section 3). To evaluate this, we track the number of steps required to reach the benchmark score on the held-out set. The target scores, derived in Dahl et al. (2023) by heavily tuning and comparing 4 different optimization algorithms, represent desirable performance on each workload.
@@
We further investigate the relationship between the optimal averaging window and training efficiency. Figure 3 reveals a consistent trend across workloads: whereas moderate weight averaging significantly reduces the number of steps required to reach the validation target, both excessively small and excessively large horizons can be suboptimal, resulting in minor speed-ups. Importantly, we find that LAWA is often highly stable and forgiving with respect to the choice of its hyperparameters, considering that for each workload a broad range of horizons are effective. This robustness is particularly evident in OGBG and FastMRI, where performance remains relatively stable across a wide range of averaging windows. We attribute this to their long training horizons, as also noted in Kasimbeg et al. (2025). Furthermore, our analysis suggests a generalizable optimal averaging horizon across workloads: an averaging window around 1% of the total training budget consistently yields optimal results. We further explore the effect of LAWA and EMA hyperparameters on training speed‐ups in Appendix B.
@@
An important question is whether these gains are to be observed only at peak tuning of an optimizer, or if they hold in generic hyperparameter settings. We investigate this circumstance by variating the top learning rate of NadamW and averaging the correspondent checkpoints. We note that LAWA and EMA maintain their efficiency gains across all considered learning rates, reliably achieving the target score faster than the baseline optimizer, as shown in Figure 4. Averaging remains effective in improving efficiency, even when applied to a suboptimal optimization algorithm. This is encouraging, as it suggests potential benefits even in the absence of a strong baseline or when searching for one is too costly. Similarly to Sanyal et al. (2023), we observe a slight advantage on OGBG with higher learning rates, but do not see a similar behavior for Criteo1TB; nevertheless on the latter WA remains effective at larger learning rates, in contrast to NadamW, which performance degrades faster. Interestingly, we notice that, as long as the baseline algorithm is able to reach the predefined target, weight averaging consistently speeds up convergence. If instead the learning rate is too big (or too small) and NadamW does not achieve the target in the maximum number of iterations, neither do the averaged checkpoints. This suggests that while averaging can significantly accelerate training, it has a limited impact on improving generalization—a point we explore further in the next section.
@@
Finally, we explore whether a more sophisticated optimizer can benefit from weight averaging. We choose Distributed Shampoo (Shi et al., 2023) as the best scoring algorithm on the AlgoPerf inaugural competition (Kasimbeg et al., 2025), which provides a significant speed-up over NadamW. We report in Figure 5 the number of steps required to reach the validation targets when training an encoder-decoder Transformer on WMT, a ViT on ImageNet and a U-Net on FastMRI. We equip Distributed Shampoo with LAWA and EMA, and observe that weight averaging consistently enhances performance, even when applied to an already highly efficient optimizer like Shampoo, reaching the targets sooner and occasionally reducing the variability of the baseline optimizer. This emphasizes that the benefits of weight averaging extend beyond compensating for slower or less effective optimizers, and indicates that averaging offers inherent advantages, regardless of the adopted optimizer.

### v3

@@ Abstract
Averaging checkpoints along the training trajectory is a simple yet powerful approach to improve the generalization performance of Machine Learning models and reduce training time. Motivated by these potential gains, and in an effort to fairly and thoroughly benchmark this technique, we present an extensive evaluation of averaging techniques in modern Deep Learning, which we perform using AlgoPerf (dahl_benchmarking_2023), a large-scale benchmark for optimization algorithms. We investigate whether weight averaging can reduce training time, improve generalization, and replace learning rate decay, as suggested by recent literature. Our evaluation across seven architectures and datasets reveals that averaging significantly accelerates training and yields considerable efficiency gains across all considered workloads, at the price of a minimal implementation and memory cost, while mildly improving generalization. Finally, we explore the relationship between averaging and learning rate annealing and show that combining the two achieves optimal performance.
@@ 1 Introduction
Training Deep Learning models is both resource-expensive and time-consuming. A principled and simple approach to speed up training, dating back to polyak_new_1990 and david_ruppert_efficient_1988, involves averaging the model’s weights across training iterations. This can either be performed online during training or post hoc by averaging checkpoints, effectively creating an ensemble of models at minimal additional cost. Previous studies have shown that weight averaging (WA) techniques can improve generalization (merity2017regularizingoptimizinglstmlanguage; Gupta2020Stochastic; kaddour_stop_2022; melis_two-tailed_2023), increase robustness (morales_ema), smooth loss landscapes (izmailov_averaging_2019), and accelerate convergence (athiwaratkun_there_2018; li_trainable_2022; sanyal_early_2023). Recent studies have also explored the connection between learning rate decaying and weight averaging (sandler_training_2023; hagele2024scaling), and used the latter to develop schedule-free algorithms (defazio_road_2024).
@@
In this work, we present the largest evaluation of averaging techniques in modern Deep Learning, which we perform using AlgoPerf (dahl_benchmarking_2023), a collection of large-scale workloads and architectures developed to provide a unified benchmark for optimization algorithms. Framing our analysis in this setting provides a carefully designed evaluation framework, enables comparisons against strong, heavily tuned baselines, and allows drawing broad and robust conclusions.
@@
    We show that averaging can significantly accelerate training dynamics across seven different models and architectures and estimate a 12% reduction in GPU-hours to train the entire AlgoPerf suite up to the validation target compared to our baseline. We find this effect to be consistent across hyperparameters and observe encouraging speed-ups even on a more sophisticated optimizer like Distributed Shampoo (shi2023distributed_shampoo).
@@
    Finally, we demonstrate how averaging checkpoints can act as a proxy for a shorter learning rate decay, but that it cannot fully replace learning rate schedules, at least within the explored variants, addressing an important question about the role of these techniques (hagele2024scaling; defazio_road_2024).
@@ 2 Related Work
The idea of averaging iterates along a stochastic optimization trajectory dates back to polyak_new_1990 and david_ruppert_efficient_1988, and is often referred to as Polyak–Ruppert averaging. It has been extensively studied in the stochastic approximation framework (polyak_acceleration_1992; bach_non-strongly-convex_2013; neu_iterate_2018; lakshminarayanan_linear_2018), and is a common technique to derive convergence guarantees in both convex and nonconvex optimization (garrigos2023handbook).
@@
Weight averaging techniques have seen extensive use in the Deep Learning community, often without explicit recognition or emphasis, despite their effectiveness. Influential works such as szegedy_2016_rethinking, vaswani2017attention, and merity2017regularizingoptimizinglstmlanguage have demonstrated their ability to enhance model performance and mitigate overfitting.
@@
The work of izmailov_averaging_2019 sparked renewed interest in weight averaging by demonstrating how averaging points along the SGD trajectory leads to wider minima and improves generalization performance. Their approach, Stochastic Weight Averaging (SWA), has since been applied to semi-supervised learning (tarvainen2018meanteachersbetterrole; athiwaratkun_there_2018), low-precision training (yang_swalp_2019), domain generalization tasks (cha_swad_2021; arpit2022ensembleaveragesimprovingmodel), and meta-optimization (li_trainable_2022).
@@
In the original formulation of SWA, a pretrained model is trained with a cyclical or constant learning rate, and multiple checkpoints are collected and later averaged. kaddour_stop_2022 proposed Latest Weight Averaging (LAWA), an online algorithm that averages the latest checkpoints in a rolling window, showing significant speed-ups on vision and language tasks. Further modifications of LAWA demonstrated notable boosts in pretraining modern decoder-only language models (sanyal_early_2023). A valuable alternative to moving window averaging techniques like LAWA is Exponential Moving Averaging (EMA) (li_switch_2024; morales_ema; arpit2022ensembleaveragesimprovingmodel). It retains similar advantages of rolling window averaging, and constitutes an indispensable technique for high-quality image synthesis models such as GANs and diffusion models (yazıcı2019avg_gan; song2021score_based_gen; karras2024nvidia_diffusion_avg).
@@
Classical stochastic smooth convex optimization rates showcase a tight link between WA and learning rate annealing, suggesting a practical interplay between these techniques (see e.g. Theorem 5.3. in garrigos2023handbook). Intuitively, averaging weights along the training trajectory reduces noise and might act as a proxy for learning rate decay. In fact, sandler_training_2023 proved the theoretical and empirical equivalence between WA and decaying learning rate for SGD. However, despite this appealing result, modern Deep Learning models are still predominantly trained with learning rate annealing (hagele2024scaling), even when maintaining an EMA of model weights (DeepSeekV3). A recent study by defazio_road_2024 specifically investigates this connection, proposing an approach that fully replaces learning rate schedules with iterate averaging and demonstrating strong performance on the same benchmark used in this analysis. Whereas defazio_road_2024 incorporates averaging directly into the optimization procedure, we explore a different flavor of averaging, where the averaged weights do not influence the updates—akin to Polyak averaging, SWA, LAWA, and EMA.
@@
Finally, a different but notable approach that leverages the benefits of averaging is model soups (wortsman_model_2022). In this case, multiple models are trained with different hyperparameter configurations and later aggregated, resulting in an ensemble with improved accuracy and robustness.
In this work, we demonstrate the benefits of weight averaging techniques on a challenging optimization benchmark (dahl_benchmarking_2023), hoping to encourage broader adoption of these methods in training large-scale Machine Learning models.
@@ 3 Experimental Setup
We explore two flavors of weight averaging discussed in Section 2: LAWA (kaddour_stop_2022) and EMA. For both approaches, we update a buffer that stores previous information about the model history. In the first case, we update a circular queue of length $L$, which stores recent checkpoints, while in the latter case, we maintain an exponential moving average of the model parameters, with coefficient $\gamma$. We update the averaging buffer every $\nu$ iterations, and explore the effect of such hyperparameter together with the value of $L$ and $\gamma$. In most scenarios, we save checkpoints of the baseline algorithm and run averaging schemes offline, but when testing collecting consecutive checkpoints, we use online versions of LAWA and EMA to avoid excessive disk storage. We discuss implementation details of the algorithms, along with their practical implications and considerations, in Appendix E.
@@
We conduct our analysis on AlgoPerf (dahl_benchmarking_2023), a large suite of Deep Learning workloads, which provides a comprehensive benchmark for testing optimization algorithms. The benchmark is composed of eight workloads, each defined by a dataset, a model architecture, and a predefined target metric on a held-out set, designed to represent optimal performance on such a workload. We note that AlgoPerf has been developed following option (A) and that significant effort has been dedicated to deriving challenging target scores. Thus, it provides an ideal setting to evaluate the impact of weight averaging techniques applied to strong optimization algorithms and compare their performance against heavily tuned baselines. We also make use of the same datasets and architectures to score algorithms by means of their generalization capabilities (when following option (B)). We consider the following workloads from the AlgoPerf suite: (i) a DLRMsmall model on Criteo 1TB dataset for click-through rate prediction; (ii) U-Net on FastMRI for medical image reconstruction; (iii) ViT on ImageNet-1k for image classification; (iv) a GNN model on OGBG for graph-based prediction; (v) a Transformer-Big on WMT for machine translation; (vi) a Conformer for speech recognition; (vii) a DeepSpeech model on LibriSpeech for speech-to-text. We exclude ImageNet‐ResNet from this analysis because, as shown in dahl_benchmarking_2023, no optimization algorithm successfully solves this task alongside the other workloads, leaving no baseline on which to build. We refer to Appendix A for more details on the workloads.
@@
We build on top of existing optimization algorithms that perform well on AlgoPerf, enhancing them with LAWA or EMA. Unless otherwise specified, we use NadamW (dozat_nadam; Loshchilov2017DecoupledWD) as our baseline and name the resulting algorithms NadamW + LAWA and NadamW + EMA, respectively. Additionally, we investigate the potential benefits of combining a higher-order optimizer, such as Distributed Shampoo (shi2023distributed_shampoo), with these averaging methods, exploring how this approach might further improve training efficiency. Given the high computational cost of the benchmark, we conduct most of the analysis using the best performing hyperparameters in dahl_benchmarking_2023 and kasimbeg2025accelerating: this gives us strong baselines, avoiding the burden of expensive hyperparameter tuning. Additionally, we ablate on the role of the learning rate in each analysis and study in more detail the consequences of changing the learning rate schedule. Given this reference algorithm, we add LAWA or EMA on top of it, tuning only the hyperparameters of the averaging scheme and leaving the other baseline hyperparameters fixed, including learning rate schedule and weight decay. Unless otherwise specified, or when explicitly ablating on it, we train using a cosine learning rate schedule (loshchilov2017sgdr).
@@
The impact of an averaging scheme largely depends on the horizon over which it is applied and on the frequency at which checkpoints are collected. A long horizon may prioritize outdated model checkpoints and generalize worse, whereas a short one may be suboptimal or result in a serious computational overhead. At the same time, updating the averaging buffers every step is only possible in the online version of LAWA and EMA, and collecting checkpoints too rarely might reduce the benefits of averaging. Previous works have collected checkpoints at each training step (morales_ema; hagele2024scaling), or the end of each epoch (kaddour_stop_2022), and investigated the optimal horizon and update frequency on a single task (sanyal_early_2023). We explore combinations and interactions of these variables across architectures and objectives, and compare LAWA with EMA.
@@ 4 Speeding up Training
We ask whether equipping a strong baseline with LAWA or EMA can reduce the number of steps required to reach a predefined validation target (option A in Section 3). To evaluate this, we track the number of steps required to reach the benchmark score on the held-out set. The target scores, derived in dahl_benchmarking_2023 by heavily tuning and comparing 4 different optimization algorithms, represent desirable performance on each workload.
@@
We further investigate the relationship between the optimal averaging window and training efficiency. Figure 3 reveals a consistent trend across workloads: whereas moderate weight averaging significantly reduces the number of steps required to reach the validation target, both excessively small and excessively large horizons can be suboptimal, resulting in minor speed-ups. Importantly, we find that LAWA is often highly stable and forgiving with respect to the choice of its hyperparameters, considering that for each workload a broad range of horizons are effective. This robustness is particularly evident in OGBG and FastMRI, where performance remains relatively stable across a wide range of averaging windows. We attribute this to their long training horizons, as also noted in kasimbeg2025accelerating. Furthermore, our analysis suggests a generalizable optimal averaging horizon across workloads: an averaging window around 1% of the total training budget consistently yields optimal results. We further explore the effect of LAWA and EMA hyperparameters on training speed‐ups in Appendix B.
@@
An important question is whether these gains are to be observed only at peak tuning of an optimizer, or if they hold in generic hyperparameter settings. We investigate this circumstance by variating the top learning rate of NadamW and averaging the correspondent checkpoints. We note that LAWA and EMA maintain their efficiency gains across all considered learning rates, reliably achieving the target score faster than the baseline optimizer, as shown in Figure 4. Averaging remains effective in improving efficiency, even when applied to a suboptimal optimization algorithm. This is encouraging, as it suggests potential benefits even in the absence of a strong baseline or when searching for one is too costly. Similarly to sanyal_early_2023, we observe a slight advantage on OGBG with higher learning rates, but do not see a similar behavior for Criteo1TB; nevertheless on the latter WA remains effective at larger learning rates, in contrast to NadamW, which performance degrades faster. Interestingly, we notice that, as long as the baseline algorithm is able to reach the predefined target, weight averaging consistently speeds up convergence. If instead the learning rate is too big (or too small) and NadamW does not achieve the target in the maximum number of iterations, neither do the averaged checkpoints. This suggests that while averaging can significantly accelerate training, it has a limited impact on improving generalization—a point we explore further in the next section.
@@
Finally, we explore whether a more sophisticated optimizer can benefit from weight averaging. We choose Distributed Shampoo (shi2023distributed_shampoo) as the best scoring algorithm on the AlgoPerf inaugural competition (kasimbeg2025accelerating), which provides a significant speed-up over NadamW. We report in Figure 5 the number of steps required to reach the validation targets when training an encoder-decoder Transformer on WMT, a ViT on ImageNet and a U-Net on FastMRI. We equip Distributed Shampoo with LAWA and EMA, and observe that weight averaging consistently enhances performance, even when applied to an already highly efficient optimizer like Shampoo, reaching the targets sooner and occasionally reducing the variability of the baseline optimizer. This emphasizes that the benefits of weight averaging extend beyond compensating for slower or less effective optimizers, and indicates that averaging offers inherent advantages, regardless of the adopted optimizer.